Gold has a face centered cubic unit cell and has the density πŸπŸ—. πŸ‘ π’ˆ π’„π’Žβˆ’πŸ‘ . Calculate the length along an edge. (𝐴𝑒 = 197 𝑔 π‘šπ‘œπ‘™ βˆ’1 )Β 


1
Expert's answer
2021-11-18T11:36:15-0500

For FCC crystal structure "n=\frac{1}{8}\cdot8+\frac{1}{2}\cdot6=4" atoms per unit cell and "V=a^3" .


"\rho=\frac{n\cdot M}{V\cdot N_A}=\frac{n\cdot M}{a^3\cdot N_A}\to a=(\frac{n\cdot M}{\rho\cdot N_A})^{1/3}="


"=(\frac{4\cdot 0.197}{19300\cdot 6.023\cdot10^{23}})^{1/3}=4.077\cdot10^{-10}\ (m)" . Answer